OpenStudy (anonymous):

Using induction to prove that 2^x < x! for all x>=4

OpenStudy (anonymous):

you know how to do a proof by induction?

OpenStudy (anonymous):

if so, after the base case this takes one step if not, then it takes the rest of the night to explain it

OpenStudy (solomonzelman):

why not for all x≥2?

OpenStudy (anonymous):

cause \[2^2>2!\]

OpenStudy (solomonzelman):

oh yes, I read the equation the other way around

OpenStudy (solomonzelman):

factorial is greater!

OpenStudy (solomonzelman):

well, you can tell that \(\large\color{#000000 }{ \displaystyle 2^4=16 }\) \(\large\color{#000000 }{ \displaystyle 4!=24 }\) multiply mutliply × 2 × larger and larger number every time every time

OpenStudy (anonymous):

is totally obvious, but you have to understand the mechanics of a proof by induction if you know what they are, this one is solved in one step as @SolomonZelman wrote above if not, then it takes for ever

OpenStudy (solomonzelman):

I recalled this logic from the very first proof of Harmonic Series Divergence - b/c that's the first time I've seen this.... I mean that: 1/5+1/6+1/7+1/8 is at least 1/8+1/8+1/8+1/8 and so on...

OpenStudy (solomonzelman):

you are multiplying times number that is at least 2, for x!